A 4-year annuity of $200 monthly payments begins 10 years from now. The required return is 10%, compounded monthly. a. What is the value of the annuity today? b. What is the value of the annuity in 3 years? c. What is the value of the annuity in 7 years? d. What is the value of the annuity in 10 years? e. What is the value of the annuity in 12 years? f. What is the value of the annuity in 15 years?
Answer #1

Answer:

Value of payment at year 10 $    7,885.63 =PV(10%/12,4*12,-200)
a Value of the annuity today $    2,913.01 =7885.63/(1+10%/12)^120
b Value of the annuity in 3 years $    3,927.26 =7885.63/(1+10%/12)^84
c Value of the annuity in 7 years $    5,849.08 =7885.63/(1+10%/12)^36
d Value of the annuity in 10 years $    7,885.63 =7885.63
e Value of the annuity in 12 years $    9,623.55 =7885.63*(1+10%/12)^24
f Value of the annuity in 15 years $ 12,974.30 =7885.63*(1+10%/12)^60